OpenStudy (anonymous):

If y=sqrt(x^2+16), then d^2y/dx^2=?

OpenStudy (owlcoffee):

okay, I will guess, that's how you denote, the second derivative of any function. That only means "derivate the derivative"... Staright forward. so, having the function and derivating it: \[y=\sqrt{x^{2}+16}\] \[\frac{ dy }{ dx }=\frac{ 1 }{ 2\sqrt{x^2+16} }(2x) \] \[\frac{ dy }{ dx }=\frac{ 2x }{ 2\sqrt{x^2+16} }\] \[\frac{ dy }{ dx }=\frac{ x }{ \sqrt{x^2+16} }\] now, you'd have to derivate it again to find d^2y/dx^2

OpenStudy (anonymous):

Okay, I got that part right. I think I just messed up something in the second derivative.

OpenStudy (owlcoffee):

show me how you did it.